OpenStudy (anonymous):

ugh idk how to do this... Simplify. Rationalize the denominator if necessary.

OpenStudy (anonymous):

\[\frac{ 4\sqrt{b} }{ \sqrt{3 - b } }\]

OpenStudy (anonymous):

first combine like terms and do radical cubic square roots and solve

OpenStudy (anonymous):

multply and divide by sqrt(3-b)

OpenStudy (anonymous):

there can never be a negative in a denominator if i recal...